# Inductive Learning of Logical Theories with LLMs: A Complexity-graded Analysis

### v1

@@ 5 Related Work
6 Conclusion
In this study we thoroughly investigate the integration of state-of-the-art formal theory induction within the context of large language models (LLMs), aiming to elucidate the extent in which LLMs can systematically perform inductive learning for theories spanning across different complexity levels. At the heart of this exploration lies the recognition of relational data’s inherent semantic depth, stemming from its symbolic representations. The empirical results presented here have indicated the ability of LLMs to address inductive learning tasks, with the largest LLMs achieving competitive results against the algorithmic SOTA with better tolerance to higher noise levels, which can be attributed to their semantic flexibility. This flexibility however has certain limitations, as we found that tested language models are more limited by their capacity of tracking long relationship chains of independent predicates than by the dependency complexity of the rule sets (disjunctiveness, recursivity).
As future work we plan to utilise larger datasets to test the scalability of the proposed approach, allowing researchers to assess its performance across a broader range of scenarios. Additionally, it is worth considering the integration of the LLM’s output as an initial input for the ILP process, potentially leveraging the strengths of both approaches to overcome their respective limitations. Another avenue, is the use of ILP middle steps, such as the bottom clause, to help LLM induce a theory.
7 Limitations
While the proposed evaluation methodology aims to cover a wide range of logic theory induction complexity, it is limited in its resolution to the categories specified by Cornelio and Thost (2021), and does not quantify other ruleset characteristics, such as workspace size or unification rate in the case of Prolog Dikovsky (1993).
@@ Acknowledgements
This work was partially funded by the Swiss National Science Foundation (SNSF) project NeuMath (200021_204617) and by the Manchester Experimental Cancer Medicine Centre and the NIHR Manchester Biomedical Research Centre.
@@ Appendix A Further theoretical background
A.1 Detailed Complexity Classes
@@
Category Disjunctive Rooted DG (DRDG). Category DRDG generalises Category RDG by allowing for alternative rules represented as children of an "OR" node. For instance:
@@
In the example, the first rule states $p7(X0,X1)$ is true if $p5(X0,X1)$ is true, indicating $p7$ depends on $p5$. The second rule states $p5(X0,X1)$ is true if $p0(X0,X1)$ is true, showing $p5$ depends on $p0$. The third rule states $p5(X0,X1)$ is true if $p8(X1,X0)$ is true, adding an alternative condition with swapped arguments. Thus, $p5$ acts as an "OR" condition in the first rule’s body and the second and third rules’ heads.
@@ Appendix G Models Output
The models in question autonomously created the head of the clause "theory," as well as the predicates "p" and "pos," which should not exist.

### v2

@@ 3. Evaluation of Hypothesis: Checking for the completeness and consistency conditions:
Conclusion
In this study we thoroughly investigate the integration of state-of-the-art formal theory induction within the context of large language models (LLMs), aiming to elucidate the extent in which LLMs can systematically perform inductive learning for theories spanning across different expressivity levels. At the heart of this exploration lies the recognition of relational data’s inherent semantic depth, stemming from its symbolic representations. The empirical results presented here have indicated the ability of LLMs to address inductive learning tasks, with the largest LLMs achieving competitive results against the algorithmic SOTA with better tolerance to higher noise levels, which can be attributed to their semantic flexibility. This flexibility however has certain limitations, as we found that tested language models are more limited by their capacity of tracking long relationship chains of independent predicates than by the dependency complexity of the rule sets (disjunctiveness, recursivity).
As future work we plan to use larger datasets to test the scalability of the proposed approach, allowing the assessment of the performance of LLMs across a broader range of domain-specific and agnostic scenarios. Additionally, we will investigate the impact of LLMs for bootstrapping the initial generation of theories.
Limitations
While the proposed evaluation methodology aims to cover a wide range of logic theory induction expressivity levels, it is limited in its resolution to the categories specified by (Cornelio and Thost 2021), and does not quantify other ruleset characteristics, such as workspace size or unification rate in the case of Prolog (Dikovsky 1993).
@@ Acknowledgements
This work was partially funded by the Swiss National Science Foundation (SNSF) project NeuMath (200021_204617) [^7] https://data.snf.ch/grants/grant/204617, by the CRUK National Biomarker Centre, and supported by the Manchester Experimental Cancer Medicine Centre and the NIHR Manchester Biomedical Research Centre.
@@ Appendix A Further theoretical background
Detailed Complexity Classes
@@
Category Disjunctive Rooted DG (DRDG). Category DRDG generalises Category RDG by allowing for alternative rules represented as children of an ”OR” node. For instance:
@@
In the example, the first rule states $p7(X0,X1)$ is true if $p5(X0,X1)$ is true, indicating $p7$ depends on $p5$. The second rule states $p5(X0,X1)$ is true if $p0(X0,X1)$ is true, showing $p5$ depends on $p0$. The third rule states $p5(X0,X1)$ is true if $p8(X1,X0)$ is true, adding an alternative condition with swapped arguments. Thus, $p5$ acts as an ”OR” condition in the first rule’s body and the second and third rules’ heads.
@@ Appendix G Models Output
The models in question autonomously created the head of the clause ”theory,” as well as the predicates ”p” and ”pos,” which should not exist.